What is the standard deviation of sampling distribution if standard deviation of population is 35 and sample size is 9?


1
Expert's answer
2022-03-10T06:57:13-0500

Let "\sigma_x" be a standard deviation of sampling distribution, "\sigma" - population standard deviation, n - sample size, then

"{\sigma}_x={\frac {{\sigma}} {\sqrt n}}={\frac {35} {\sqrt 9}}={\frac {35} 3}\approx11.67"